OpenStudy (anonymous):

Omar the Rope Maker decided to make a rope long enough to stretch around the Earth (the circumference of which is 40000 kilometers). Having completed the task, he discovered that he had actually made the rope 12 metres too long. Now wanting to cut the rope (it was somewhat of a record length), he decided to put the ends of the rope together and have all his relatives help him hold the rope and equal distance off the ground all the way around the earth. Which of these animals could pass under the rope: (a) ant (b) snake (c) Omar (d) adult elephant (e) blue whale?

OpenStudy (anonymous):

We'd want to find the difference between the radius of the rope above the earth and the radius of the earth. C=2*pi*r r=C/(2*pi) Converting km to m as well 40000012m/(2*pi)-40000000/(2*pi) is approximately 1.91 meters The rope is about 1.91 m above the Earth's surface. Not sure of the height of all those animals, so you'd have to research that.

OpenStudy (anonymous):

A solution using Mathematica is attached.